# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 10 Reinforcement Learning from Human Feedback
Every chapter so far assumes access to a scalar reward signal. Dynamic programming (Chapter 3) requires the full reward function $r(s,a)$. Model-free RL (Chapter 2) relaxes the model requirement but still observes a numerical reward $r_{t}$ after each transition. Q-learning, temporal difference methods, policy gradient, all update toward a scalar signal. The Bellman optimality equation $V^{*}(s)=\max_{a}\{r(s,a)+\gamma\sum_{s^{\prime}}P(s^{\prime}|s,a)V^{*}(s^{\prime})\}$ presupposes that $r(s,a)$ is either known or observable. When it is neither, the entire DP/RL machinery cannot be applied directly.
In many domains, scalar rewards are unavailable but ordinal preferences over trajectories are easy to elicit. Text generation is the canonical example: a human evaluator cannot assign a meaningful numerical score to a paragraph of text, but can reliably say “response A is better than response B” when shown two candidate completions. Similarly, in robotics, a human can compare two demonstrated behaviors without specifying a cost function. The raw data in RLHF is a set of trajectory pairs (or trajectory segment pairs) with binary preference labels indicating which trajectory the evaluator preferred. RLHF uses these ordinal comparisons to learn a proxy reward function $r_{\theta}$, which then serves as the scalar signal for standard RL optimization. This is not an inverse problem in the IRL sense; the goal is not to rationalize observed behavior or recover the agent’s own reward, but rather a two-stage forward problem: first learn a reward from human judgments, then solve the resulting MDP using the RL tools from earlier chapters.
When no one can write down the reward function, human preferences provide an alternative training signal. Reinforcement Learning from Human Feedback (RLHF) learns a reward function by eliciting pairwise preference comparisons from a human evaluator over machine-generated outputs, rather than inverting an observed policy as in structural estimation or IRL. Christiano et al. (2017) demonstrated that this approach could train agents without an explicit reward function. RLHF has since become the predominant method for aligning large language models.
@@ 10.1 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a direct application of discrete choice theory and is equivalent to a binary logit model, a special case of the multinomial logit framework (McFadden, 1974a) that underpins DDC estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy, so preferences are over trajectories rather than single actions. The preference loss in Equation (74) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 10.2 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT) on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third and final step, the policy is optimized using reinforcement learning. The $\pi^{SFT}$ policy, now denoted $\pi_{\phi}$, is fine-tuned to maximize the expected reward signal from the frozen reward model $r_{\theta}$. In practice, this optimization uses PPO (Chapter 2, Section 5), the same clipped policy gradient method that stabilizes updates via a trust region. The standard policy gradient machinery from Chapter 2 applies directly, with one twist: the reward function is itself an estimate, introducing approximation error on top of the usual variance in gradient estimates. A critical component of this optimization is the inclusion of a regularization term that penalizes the Kullback-Leibler (KL) divergence between the learned policy $\pi_{\phi}$ and the initial reference policy $\pi^{SFT}$. The objective function for this final stage is:
@@
This formulation directly parallels Bayes’ rule, where the posterior is proportional to the prior times the likelihood. The reward function, scaled by $\lambda_{KL}$, defines the log-likelihood of the preference evidence. From this perspective, the KL-regularized RL objective $J(\phi)$ is no longer arbitrary. It is mathematically equivalent, up to an additive constant, to the Evidence Lower Bound (ELBO) for this Bayesian inference problem. Consequently, maximizing the RLHF objective via PPO is a form of variational inference, where the aim is to find the policy $\pi_{\phi}$ that minimizes the KL divergence to the ideal posterior $\pi^{*}$. This insight reframes RLHF from a reward-maximization task into a probabilistic inference problem, providing a theoretical justification for the KL penalty as a structural component of the inference problem rather than an ad-hoc regularizer.
Despite this closed-form characterization, the three-stage RLHF pipeline is complex to implement, requiring training multiple large models and a computationally expensive reinforcement learning loop that involves sampling from the policy. Building on the analytical relationship between the optimal policy and the reward function, Direct Preference Optimization (DPO) was introduced by Rafailov et al. (2023) as a simpler, RL-free method. DPO eliminates the RL optimization loop entirely, bypassing the policy gradient and value function machinery of Chapters 2 and 3. Where standard RLHF decomposes the problem into “learn reward from trajectory preferences” then “solve the resulting MDP,” DPO solves for the policy directly from the same preference data over trajectory pairs, collapsing two stages into one supervised learning objective. DPO bypasses the explicit reward modeling step by reparameterizing the reward function directly in terms of the optimal policy $\pi^{*}$ and the reference policy $\pi^{SFT}$, as in Equation (77):
@@
This objective can be optimized directly on the policy parameters $\phi$ using standard supervised learning techniques with a static preference dataset. The gradient of this loss intuitively increases the likelihood of preferred responses $y_{w}$ while decreasing the likelihood of dispreferred responses $y_{l}$, relative to the reference policy. By recasting the constrained optimization problem as a simple likelihood maximization, DPO offers a more stable and computationally efficient alternative to the multi-stage RLHF pipeline, eliminating the need for an explicitly trained reward model and the complexities of RL-based policy optimization.
@@ 10.3 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.135135135The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the “alignment tax” (Ouyang et al., 2022).136136136The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
10.4 Simulation Study: Learning from Preferences on a Gridworld
The gridworld is a $10\times 10$ stochastic grid with slip probability $p=0.1$ and discount factor $\gamma=0.99$, structurally identical to the environments in Chapters 3a and 4. The agent starts at $(0,0)$ and must reach the goal at $(9,9)$, receiving r=0.1r=-0.1 per step, $r=+10$ for reaching the goal, r=2r=-2 for entering hazard cells at $(3,3)$ and $(6,7)$, and r=5r=-5 for entering a trap at $(5,2)$. The true reward function is linear in four indicator features, $r(s,a,s^{\prime})=\phi_{0}+\phi_{1}\mathbbm{1}\{s^{\prime}=\text{goal}\}+\phi_{2}\mathbbm{1}\{s^{\prime}\in\text{hazards}\}+\phi_{3}\mathbbm{1}\{s^{\prime}\in\text{traps}\}$, with $\phi=(-0.1,10.1,-1.9,-4.9)$. Dynamic programming gives $V^{*}(s_{0})=6.55$; the optimal policy avoids all hazards and traps. Figure 7 shows the environment with $V^{*}(s)$ and $\pi^{*}$.
Figure 7: The $10\times 10$ stochastic gridworld. Cell shading shows $V^{*}(s)$ (viridis scale); arrows indicate the optimal policy $\pi^{*}$. S marks the start, G the goal, H the hazard cells, and T the trap. The optimal policy routes right and down toward the goal while steering around hazards and traps.
Preference data is generated by rolling out a uniform random policy (each action equally likely) from random non-goal start states. Each rollout produces a trajectory segment of $L=25$ steps. For each of $K$ comparisons, two independent segments are generated; the segment with higher cumulative reward under the true (unobserved) reward function is labeled as preferred via the Bradley-Terry model. This produces $K$ binary comparisons. All six methods receive identical preference data per seed; the agent never observes per-transition rewards, only the binary preference label per pair.
Five methods are compared across $K\in\{25,50,100,200,500,1000,2000,5000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP (8 inputs encoding normalized row and column positions plus one-hot action, 64 hidden units, $\sim$4,800 parameters) trained by Bradley-Terry MLE on the $K$ segment pairs. For each segment, the model computes per-transition rewards $r_{\theta}(s_{t},a_{t},s_{t+1})$ and sums them to obtain a segment score; the preferred segment should score higher under the logistic loss from Equation (74). After 200 training epochs (learning rate $10^{-3}$, batch size 64), the $100\times 4$ reward table is extracted and the resulting MDP solved by value iteration. The correctly specified structural model parameterizes reward as $r=\phi_{0}+\phi_{1}\mathbbm{1}\{\text{goal}\}+\phi_{2}\mathbbm{1}\{\text{hazard}\}+\phi_{3}\mathbbm{1}\{\text{trap}\}$ (4 parameters matching the true form); $\phi$ is fit by L-BFGS-B maximization of the same Bradley-Terry log-likelihood, and the resulting MDP solved by value iteration. The misspecified variant uses only $\phi_{0}$ and $\phi_{1}$, ignoring hazards and traps. Tabular Q-learning (20,000 episodes, $\varepsilon=0.1$, $\alpha=0.1$) and exact DP provide scalar-reward baselines that observe per-transition rewards.
Table 15: Sample complexity of preference-based policy learning on a stochastic $10\times 10$ gridworld (30 seeds). The neural network reward model reaches 99% of DP-optimal at K=1, [^0] K=1{,}000. The correctly specified structural model reaches 100% at K=5, [^0] K=5{,}000 but has high variance at intermediate $K$. The misspecified model plateaus near 98% while routing through hazard cells.
Method $K$ Return % of DP
DP Oracle $6.52\pm 0.02$ $100.0$
Q-Learning $6.50\pm 0.01$ $99.7$
NN RLHF $25$ $6.28\pm 1.29$ 96.3-96.3
Correct $25$ $4.89\pm 0.90$ $75.0$
Misspecified $25$ $4.79\pm 0.90$ $73.5$
NN RLHF $50$ $3.37\pm 1.43$ 51.6-51.6
Correct $50$ $6.00\pm 0.54$ $92.0$
Misspecified $50$ $5.89\pm 0.54$ $90.2$
NN RLHF $100$ $3.06\pm 1.19$ $47.0$
Correct $100$ $4.35\pm 1.02$ $66.6$
Misspecified $100$ $4.81\pm 0.90$ $73.8$
NN RLHF $200$ $4.75\pm 0.90$ $72.8$
Correct $200$ $4.90\pm 0.90$ $75.2$
Misspecified $200$ $4.82\pm 0.90$ $73.8$
NN RLHF $500$ $6.21\pm 0.27$ $95.1$
Correct $500$ $4.89\pm 0.90$ $75.0$
Misspecified $500$ $4.78\pm 0.90$ $73.3$
NN RLHF $1000$ $6.45\pm 0.04$ $98.9$
Correct $1000$ $5.46\pm 0.75$ $83.6$
Misspecified $1000$ $5.89\pm 0.54$ $90.3$
NN RLHF $2000$ $6.52\pm 0.01$ $99.9$
Correct $2000$ $6.01\pm 0.54$ $92.1$
Misspecified $2000$ $6.44\pm 0.02$ $98.7$
NN RLHF $5000$ $6.48\pm 0.01$ $99.4$
Correct $5000$ $6.55\pm 0.00$ $100.4$
Misspecified $5000$ $6.42\pm 0.02$ $98.4$
Table 15 and Figure 8 report the main results. The neural network reward model converges fastest, rising from negative returns at $K=25$ to 95% of DP-optimal at $K=500$ and 99% at K=1, [^0] K=1{,}000. The correctly specified structural model dominates at very low $K$ (75% at $K=25$ versus $96\%$ for the neural net) but exhibits high variance at intermediate $K$ (standard errors of $\pm 0.9$), because hazard and trap cells constitute only 3 of 99 non-terminal states and random segments rarely visit them, leaving $\phi_{2}$ and $\phi_{3}$ poorly identified.137137137At $K=100$, the estimated trap penalty $\hat{\phi}_{3}=-18.6$ overshoots the true value of 4.9-4.9 by a factor of four, causing the policy to take extreme detours. At K=5, [^0] K=5{,}000, the structural model achieves 100.4% of DP return with parameter estimates converging toward their true values. The misspecified model converges to 98.4%, a seemingly small gap, but its policy consistently routes through hazard cells (2.1 visits on average), confirming qualitative misspecification.138138138The modest return gap arises because hazard penalties (2-2 per visit, $\sim$2 visits per episode) are small relative to total return ($\sim$6.5). Larger penalties or more hazards would widen it. Q-learning converges to 99.7% of DP using 458,000 scalar reward observations, three orders of magnitude more signal than the neural net requires at K=1, [^0] K=1{,}000. The flexible neural net outperforms the correctly specified structural model at intermediate sample sizes ($K=500$ through K=2, [^0] K=2{,}000) because its 8-dimensional input representation extracts reward information from every transition, not just the rare visits to 3 special cells. Even mild misspecification (omitting 2 of 4 reward features) produces qualitatively wrong policies despite quantitatively similar returns.
Table 16: Diagnostics at K=5, [^0] K=5{,}000 (single seed). Policy agreement measures the fraction of non-goal states where each method’s greedy policy matches $\pi^{*}$. Reward MAE is the mean absolute error of the estimated expected reward table against the true table. $V^{\pi}$ correlation is the Pearson correlation between the true value of the learned policy (evaluated under the true reward) and $V^{*}$. Hazard visits counts states where the deterministic policy routes into hazard or trap cells.
Method Policy agreement (%) Reward MAE $V^{\pi}$ corr. with $V^{*}$ Hazard visits
NN RLHF $77.8$ $0.766$ $0.836$ 11
Correct $100.0$ $0.081$ $1.000$ 0
Misspecified $77.8$ $0.190$ $0.809$ 22
Table 16 provides a different view of the K=5, [^0] K=5{,}000 results, showing what each method actually learned rather than just policy returns. The correctly specified structural model achieves 100% policy agreement with $\pi^{*}$ and value function correlation 1.000, recovering the true reward to within MAE 0.081. The neural network model, despite achieving 99% of DP return in the main experiment, agrees with $\pi^{*}$ on only 78% of states; the disagreements occur at states far from the typical trajectory where the precise action choice has little impact on expected return. The misspecified model matches the neural net in policy agreement (78%) but for a different reason; it routes through hazard cells that the optimal policy avoids, producing a qualitatively distinct trajectory distribution with value function correlation 0.809.
Table 17: Online versus offline data collection for the neural network RLHF model at K=1, [^0] K=1{,}000 (20 seeds). Online collection is unstable; catastrophic failures on approximately half of seeds produce a negative mean return.
Collection Return $p$-value
Online (4 rounds) $2.59\pm 1.81$
Offline $5.82\pm 0.66$ $0.9999$
Table 17 reports the online versus offline ablation for the neural network model at K=1, [^0] K=1{,}000. The online procedure divides $K$ into four rounds of 250 comparisons each, collecting subsequent rounds from the current learned policy rather than the random policy. Offline collection ($5.82\pm 0.66$) substantially outperforms online ($2.59\pm 1.81$, $p>0.999$). A poor first-round reward model, trained on only 250 comparisons, produces a degenerate policy whose concentrated state visitation degrades subsequent rounds through a compounding feedback loop. In this gridworld, the random policy already provides near-uniform state coverage that targeted collection can only narrow. The finding contrasts with Christiano et al. (2017), where online collection helped in high-dimensional Atari and MuJoCo environments because random exploration fails to cover the relevant state space; in small, well-covered environments, the instability risk of online collection outweighs its potential benefit.
Figure 8: Policy return versus number of preference comparisons. The neural network RLHF model ($\sim$4,800 parameters) converges fastest to near-optimal return. The structural model (4 parameters) has high variance at intermediate $K$ but achieves exact recovery at K=5, [^0] K=5{,}000.
11 Offline Reinforcement Learning and Causal Inference

### 15 Mar 2026 v2

@@ 12 Reinforcement Learning from Human Feedback
Every chapter so far assumes access to a scalar reward signal: dynamic programming requires $r(s,a)$, model-free RL observes $r_{t}$ after each transition, and the Bellman optimality equation presupposes that rewards are known or observable. When they are neither, the DP/RL machinery cannot be applied directly.
In many domains, scalar rewards are unavailable but ordinal preferences over trajectories are easy to elicit. A human evaluator cannot assign a meaningful numerical score to a paragraph of text, but can reliably say “response A is better than response B.” The raw data is a set of trajectory pairs with binary preference labels. RLHF uses these ordinal comparisons to learn a proxy reward function $r_{\theta}$, which then serves as the scalar signal for standard RL optimization. This is not an inverse problem in the IRL sense; the goal is not to rationalize observed behavior, but rather a two-stage forward problem in which the analyst first learns a reward from human judgments and then solves the resulting MDP. Christiano et al. (2017) demonstrated that this approach could train agents without an explicit reward function. RLHF has since become the predominant method for aligning large language models.
@@ 12.1 Learning Rewards from Preferences
The canonical RLHF framework is built on a formal model of human preference. The observed data consists of tuples $(s,y_{w},y_{l})$, where $s$ is a context, and $y_{w}$ and $y_{l}$ are two outputs, with $y_{w}$ being the “winner” preferred by a human. Assuming preferences follow a latent utility model, the Bradley-Terry model (Bradley and Terry, 1952) gives the probability that $y_{w}$ is preferred: $P(y_{w}\succ y_{l}|s)=\sigma(r_{\theta}(s,y_{w})-r_{\theta}(s,y_{l}))$, where $r_{\theta}:\mathcal{S}\times\mathcal{Y}\to\mathbb{R}$ is a learned reward model parameterized by $\theta$, trained to approximate human preferences (not the ground-truth reward, which is unobserved), and $\sigma(\cdot)$ is the logistic function. This formulation is a binary logit model (Section 2, Equation 1).187187187Iskhakov et al. (2020) discuss the contrasts and synergies between machine learning and structural econometrics, including the shared reliance on logit-based choice models that underlies both RLHF and dynamic discrete choice estimation. The reward model parameters $\theta$ are estimated by minimizing the negative log-likelihood of the observed human choices:
@@
In the LLM setting, the “outputs” $y_{w}$ and $y_{l}$ are token sequences, i.e., trajectories of the autoregressive policy188188188An autoregressive language model generates text token by token: at each step it outputs a distribution over the vocabulary conditioned on all preceding tokens, then samples the next token; the full response is therefore a trajectory in token space and the model acts as a sequential policy over a vocabulary-sized action set., so preferences are over trajectories rather than single actions. The preference loss in Equation (96) is MLE of a choice model where the “alternatives” are trajectory segments and the “choice” is the human-preferred one.
@@ 12.2 The RLHF Pipeline and Direct Optimization
The learned $r_{\theta}$ then serves as a proxy objective for policy optimization. Building on the reward-learning and RL fine-tuning framework of Ziegler et al. (2019) and Stiennon et al. (2020), the canonical three-stage pipeline was formalized by Ouyang et al. (2022). First, a base pretrained model is initialized via supervised fine-tuning (SFT)189189189Pretraining optimizes a language model to predict the next token across a massive text corpus, producing broad linguistic knowledge with no behavioral objective. Supervised fine-tuning (SFT) continues training on a small curated dataset of (prompt, ideal-response) pairs to specialize the model toward the desired task and establish the reference policy $\pi^{SFT}$ from which the KL penalty is measured. on a small dataset of high-quality demonstrations, yielding an initial policy $\pi^{SFT}$. This grounds the model in the desired style and format. The second step is the reward model training as described, using preference data generated from this $\pi^{SFT}$.
In the third step, the SFT policy $\pi_{\phi}$ is fine-tuned via PPO (Section 5.5) to maximize the frozen reward model $r_{\theta}$, with a KL-divergence penalty preventing the policy from drifting into regions where $r_{\theta}$ is unreliable. The objective is
@@
The reward function scaled by $\lambda_{KL}$ defines the log-likelihood, so the KL-regularized objective $J(\phi)$ is equivalent (up to an additive constant) to the Evidence Lower Bound (ELBO) for this Bayesian inference problem. Maximizing the RLHF objective via PPO is therefore variational inference: finding the policy $\pi_{\phi}$ that minimizes KL divergence to $\pi^{*}$. This reframes the KL penalty as a structural component of the inference problem rather than an ad-hoc regularizer.
Despite this closed-form characterization, the three-stage RLHF pipeline is complex to implement, requiring training multiple large models and a computationally expensive RL loop. Direct Preference Optimization (DPO), introduced by Rafailov et al. (2023), collapses the pipeline into a single supervised learning objective by reparameterizing the reward function in terms of $\pi^{*}$ and $\pi^{SFT}$, as in Equation (99):
@@
This objective is optimized on $\phi$ using standard supervised learning with a static preference dataset. The gradient increases the likelihood of preferred responses $y_{w}$ and decreases the likelihood of dispreferred responses $y_{l}$, relative to the reference policy.
Figure 21: RLHF versus DPO pipelines. Top row: the three-stage RLHF pipeline trains a reward model from human preferences, then uses PPO to fine-tune the policy with a KL penalty. Bottom row: DPO collapses the pipeline into a single supervised learning objective over preference pairs, eliminating the explicit reward model (ghosted box).
@@ 12.3 Recent Developments
While RLHF and its successors have proven effective at aligning model behavior with human preferences, several open issues remain. The framework does not aim to solve the identification problem in the strict econometric sense; the learned reward model $r_{\theta}$ (whether explicit or implicit in DPO) is a proxy for preference, not necessarily the uniquely identified “true” utility function required for welfare analysis.191191191The identification problem in preference learning mirrors that in discrete choice. The reward function is identified only up to an additive constant and requires a location normalization. Furthermore, the notion of “human feedback” is a significant simplification, as the preferences being optimized are those of a small, non-representative group of paid labelers, raising important questions about whose values are being embedded in these systems. Finally, the fine-tuning process can lead to degraded performance on standard academic benchmarks, a phenomenon called the alignment tax (Ouyang et al., 2022).192192192The alignment tax, as described by Ouyang et al. (2022), refers to performance regressions on public NLP benchmarks (SQuAD, DROP, HellaSwag) that result from RLHF fine-tuning. Mitigating these challenges while scaling alignment beyond the limits of human data collection remains a key frontier for the field.
12.4 Simulation Study: Preference Learning in Job Search
A worker searches for jobs in a labor market with compensating differentials, following a McCall (1970)-style search model. Each job is characterized by a wage $w\in\{20,28,38,50,65,82,100,125\}$ (thousands) and an amenity level $z\in\{0,1,\ldots,6\}$ capturing commute quality, flexibility, and job security. The state space has 112 states: 56 searching states in which the worker observes a pending offer $(w_{i},z_{j})$ and decides to accept or reject, plus 56 employed states $(w_{i},z_{j})$ in which the worker decides to stay or quit. The offer distribution exhibits compensating differentials, with wage rank and amenity rank negatively correlated ($\rho=-0.74$): high-wage offers cluster with low amenities and vice versa. The worker’s true per-period utility is $u(w,z)=\alpha\log(w)+(1-\alpha)z$ with $\alpha=0.6$, but this function is unobserved; the worker can only compare career trajectories (“I prefer path A to path B”), exactly as in stated-preference surveys in labor economics. While searching, the worker receives the unemployment benefit $u_{b}=\alpha\log(b)$ where $b=28$. Layoffs occur with probability $p=0.05$ per period, and the discount factor is $\gamma=0.95$. Dynamic programming gives $V^{*}(s_{0})=74.13$; the optimal policy accepts 25 of 56 offer types and stays employed at 25 of 56 job types. Preference data is generated by rolling out a uniform random policy from a random searching state. Each rollout produces a career segment of $L=15$ periods recording states and actions. For each of $K$ comparisons, two independent career segments are generated; the segment with higher cumulative discounted utility under the true (unobserved) utility function is labeled as preferred via the Bradley-Terry model. Figure 22 shows the optimal accept/reject and stay/quit boundaries.
Figure 22: Left: the optimal accept/reject boundary for searching states in wage-amenity space. Right: the optimal stay/quit boundary for employed states. Both boundaries cut diagonally, reflecting the tradeoff between wage and amenity quality under compensating differentials.
Six methods are compared across $K\in\{25,50,100,200,500,1{,}000,2{,}000,5{,}000\}$, averaged over 30 seeds. The neural network RLHF reward model is a two-layer MLP trained by Bradley-Terry MLE on the $K$ segment pairs; per-transition rewards are discount-weighted and summed to obtain a segment score, and the resulting 112-state reward table is solved by value iteration.193193193The neural network has 4 inputs (normalized log-wage, normalized amenity, employment indicator, action), 32 hidden units per layer, and $\sim$1,200 parameters. The logistic loss from Equation (96) is applied to discount-weighted segment reward sums. The correctly specified structural model parameterizes utility as $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})z$, estimating the single parameter $\hat{\alpha}$ via Bradley-Terry MLE, then solves the MDP. The misspecified model uses $\hat{u}=\hat{\alpha}\log(w)+(1-\hat{\alpha})\bar{z}$, where $\bar{z}=3$ is the mean amenity level; this model ignores amenity variation across jobs, treating all amenities as identical. DPO trains a tabular softmax policy directly from trajectory comparisons, bypassing reward modeling entirely.194194194DPO uses 112 logit parameters $\phi_{s}$, one per state, trained via the DPO loss (Equation 100) with Adam optimization over a sweep of $\lambda_{KL}\in\{0.01,0.05,0.1,0.5,1.0,5.0\}$, selecting the $\lambda_{KL}$ that minimizes training loss. The reference policy is uniform: $\pi^{SFT}(a|s)=0.5$. To match the LLM setup where both completions condition on the same prompt, DPO comparison pairs start from the same initial state. Tabular Q-learning (10,00010{,}000 episodes, $\varepsilon=0.15$, learning rate $0.1$) and exact DP provide scalar-reward baselines. All four preference methods receive identical comparison data per seed; DPO receives a same-state variant generated from the same seed.
Table 23: Sample complexity of preference-based policy learning on a McCall-style job search MDP with compensating differentials (30 seeds, $L=15$). Six methods spanning four preference-learning approaches (structural, neural network, DPO) and two scalar-reward baselines (DP, Q-learning).
Method $K$ $V^{\pi}(s_{0})$ % of DP
DP Oracle $74.13$ $100.0$
Q-Learning $73.60\pm 0.05$ $99.3$
NN RLHF $25$ $73.05\pm 0.13$ $98.5$
DPO $25$ $66.39\pm 0.34$ $89.6$
Correct $25$ $74.03\pm 0.03$ $99.9$
Misspecified $25$ $67.34\pm 0.38$ $90.8$
NN RLHF $50$ $73.31\pm 0.08$ $98.9$
DPO $50$ $68.02\pm 0.23$ $91.8$
Correct $50$ $74.07\pm 0.02$ $99.9$
Misspecified $50$ $67.12\pm 0.33$ $90.6$
NN RLHF $100$ $73.60\pm 0.06$ $99.3$
DPO $100$ $69.25\pm 0.18$ $93.4$
Correct $100$ $74.10\pm 0.01$ $100.0$
Misspecified $100$ $67.24\pm 0.28$ $90.7$
NN RLHF $200$ $73.85\pm 0.05$ $99.6$
DPO $200$ $70.16\pm 0.16$ $94.6$
Correct $200$ $74.12\pm 0.00$ $100.0$
Misspecified $200$ $67.54\pm 0.17$ $91.1$
NN RLHF $500$ $73.99\pm 0.02$ $99.8$
DPO $500$ $70.60\pm 0.10$ $95.2$
Correct $500$ $74.13\pm 0.00$ $100.0$
Misspecified $500$ $67.55\pm 0.15$ $91.1$
NN RLHF $1000$ $73.95\pm 0.03$ $99.8$
DPO $1000$ $70.59\pm 0.09$ $95.2$
Correct $1000$ $74.12\pm 0.00$ $100.0$
Misspecified $1000$ $67.52\pm 0.12$ $91.1$
NN RLHF $2000$ $74.04\pm 0.02$ $99.9$
DPO $2000$ $70.41\pm 0.07$ $95.0$
Correct $2000$ $74.13\pm 0.00$ $100.0$
Misspecified $2000$ $67.41\pm 0.11$ $90.9$
NN RLHF $5000$ $74.09\pm 0.01$ $99.9$
DPO $5000$ $70.14\pm 0.07$ $94.6$
Correct $5000$ $74.13\pm 0.00$ $100.0$
Misspecified $5000$ $67.46\pm 0.11$ $91.0$
Figure 23: Policy value $V^{\pi}(s_{0})$ versus number of preference comparisons $K$ for all six methods.
Table 23 and Figure 23 report the main results. Three findings stand out. First, the correctly specified structural model dominates: even at $K=25$ it achieves 99.9% of DP-optimal, illustrating the power of correct specification when the model has a single free parameter. Second, the neural network converges more slowly but reaches 99.9% by K=5, [^0] K=5{,}000, reflecting the higher sample complexity of a flexible model relative to a one-parameter structural specification. Third, DPO plateaus at approximately 95% by $K=500$ and does not improve further.195195195DPO learns only from $(s,a)$ pairs visited in training trajectories generated by the random behavioral policy. It cannot propagate value to undervisited states the way value iteration does after learning a reward model, so states poorly covered by the behavioral policy remain suboptimal regardless of $K$. This plateau is qualitatively different from the gridworld failure ($118\%$): DPO recovers a reasonable policy, but the gap does not close with additional data.196196196DPO fails catastrophically in gridworld because transitions are stochastic (10% slip probability) and rewards are transition-dependent. The same $(s,a)$ pair yields different rewards depending on whether the agent slipped, so the DPO loss conflates policy quality with transition luck. In the job search model, accept/reject deterministically changes employment status, and only the 5% layoff probability introduces stochastic transitions. The misspecified constant-amenity model plateaus at 91% regardless of $K$, because additional preference data cannot correct the omitted variable.
Table 24: Diagnostics at K=5, [^0] K=5{,}000 (single seed): policy agreement with $\pi^{*}$, value-function correlation, and mean accepted wage and amenity for each method.
Method Policy agree. (%) $V^{\pi}$ corr. Mean amenity Mean wage $\hat{\alpha}$
NN RLHF $96.4$ $1.000$ $4.67$ $73$
DPO $57.1$ $0.777$ $3.38$ $63$
Correct $100.0$ $1.000$ $4.84$ $71$ $0.597$
Misspecified $50.0$ $0.889$ $3.00$ $70$
Optimal $100.0$ $1.000$ $4.84$ $71$
Table 24 provides state-level diagnostics at K=5, [^0] K=5{,}000. The structural model and neural network achieve near-perfect policy agreement with $\pi^{*}$ (100% and 96% respectively). DPO agrees on only 57% of states with value-function correlation 0.78, systematically underselecting on both wage and amenity dimensions.197197197DPO’s mean accepted amenity of 3.4 parallels the misspecified structural model’s 3.0, though the mechanisms differ: the misspecified model ignores amenity variation by construction, while DPO underweights amenities because the random behavioral policy underrepresents high-amenity employed states in the training data. The misspecified model agrees on 50%, with disagreements concentrated where amenity variation matters.198198198An online versus offline ablation for the neural network at K=1, [^0] K=1{,}000 (20 seeds) shows comparable performance: online $73.92\pm 0.05$, offline $73.99\pm 0.03$ ($p=0.09$). The random behavioral policy already provides diverse career trajectories covering the full wage-amenity space.
Table 25: Segment length ablation at K=2, [^0] K=2{,}000 (20 seeds). Policy value $V^{\pi}(s_{0})$ for the neural network and DPO across segment lengths $L\in\{1,3,5,10,15,25\}$.
$L$ NN RLHF $V^{\pi}(s_{0})$ % of DP DPO $V^{\pi}(s_{0})$ % of DP
11 $56.62\pm 2.28$ $76.4$ $49.66\pm 0.25$ $67.0$
33 $73.89\pm 0.04$ $99.7$ $69.31\pm 0.11$ $93.5$
55 $74.01\pm 0.02$ $99.8$ $70.04\pm 0.09$ $94.5$
$10$ $74.02\pm 0.03$ $99.9$ $70.29\pm 0.08$ $94.8$
$15$ $74.05\pm 0.02$ $99.9$ $70.57\pm 0.07$ $95.2$
$30$ $74.05\pm 0.02$ $99.9$ $70.53\pm 0.13$ $95.1$
Figure 24: Policy value $V^{\pi}(s_{0})$ versus segment length $L$ at K=2, [^0] K=2{,}000 for the neural network and DPO.
Table 25 and Figure 24 report the segment length ablation at K=2, [^0] K=2{,}000. At $L=1$, both methods perform poorly because single-step comparisons carry minimal information about long-run value. The neural network recovers rapidly (by $L=3$) because the reward model aggregates per-transition estimates over longer segments and value iteration propagates them through the full transition structure; DPO improves monotonically but plateaus at its $\sim$95% ceiling, as it must learn the policy directly from trajectory comparisons without access to the transition model.
Two-stage RLHF separates preference estimation (a static econometric problem) from dynamic programming (which exploits the known transition model); DPO conflates the two and forfeits the transition structure that economists typically have access to.199199199This advantage is specific to settings where the transition model is known or estimable; in domains without a tractable transition model, DPO’s single-stage approach avoids compounding errors from reward model estimation.
13 Reinforcement Learning and Causal Inference